OpenStudy (anonymous):

How would I solve a question like the square root of eight over nine. Or four times the square root of eight divided by the square root of 2.

OpenStudy (anonymous):

1.) \[\sqrt{\frac{8}{9}}=\frac{\sqrt{8}}{\sqrt{9}}=\frac{2\sqrt{2}}{3}\]

OpenStudy (anonymous):

\[\frac{4\sqrt{8}}{\sqrt{2}}=\frac{4\sqrt{2}^3}{\sqrt{2}}=4\sqrt{2}^2=4*2\]